solve the system: 4x+3y-11z=8
x+y-3z=3

Answer by venugopalramana(3286)

solve the system: 4x+3y-11z=8.................1
x+y-3z=3.......................2
there are 2 eqns. and 3 unknowns.hnce we shall get infinite solutions.
eqn.1 - 3*eqn.2
4x+3y-11z-3x-3y+9z=8-9=-1
x-2z = -1...........................3
hence take any value for x say 1
1-2z=-1
z=1
put in eqn.2
1+y-3=3
y = 5
x=1,y=5,z=1 is a solution set.
similarly give any value to x and find y and z using the eqns.3 and 2
in general we can say that
x=x(any value);y=(x+9)/2 ;z=(x+1)/2 is a solution set